OpenStudy (anonymous):

The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression 80*2^-t/28 grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression represent? How would i do this equation?

OpenStudy (lgbasallote):

you substitute for t..

OpenStudy (anonymous):

for all of them? 0 , 28, n 56?

OpenStudy (lgbasallote):

let x be the final population...it says the equation is \[x = 80 \times 2^{-t/28}\] so yeah substitute t = 0, 28 and 56 i.e. when t = 0 \[x = 80 \times 2^0 \implies 80 \times 1 \implies 80\] like that

OpenStudy (anonymous):

then i would do the same for 28 and 56?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

good news for you... you dont need a calculator for that =))

OpenStudy (anonymous):

okay thanks.. can i post my answer and you'll check for me to see if their right? lols.

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

is it like this x=80*2^-0/28? or without the /28?

OpenStudy (lgbasallote):

well...0/28 is just same with 0 without /28

OpenStudy (lgbasallote):

it still becomes 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay so for 80*2^-28/28 , x=80 80*2^-56/28 = 80 Is this right?

OpenStudy (lgbasallote):

yup.. btw just so you know...that was the example i did above hehe lol

OpenStudy (lgbasallote):

^i meant the 0

OpenStudy (lgbasallote):

wait...why is everything equal to 80?

OpenStudy (anonymous):

oh wait the 80*2^-28/28 i got x=1/16777216

OpenStudy (anonymous):

and uh i dont know why everything equals to 80..

OpenStudy (lgbasallote):

you do 2^-28/28 before multiplying by 80

OpenStudy (anonymous):

how would i put 2^28/28 into a calculator?

OpenStudy (lgbasallote):

\[\Large 80 \times 2^{-28/28} \implies 80\times 2^{-1} \implies 80 \times \frac 12\]

OpenStudy (lgbasallote):

that was why i was saying you wont need a calculator

OpenStudy (anonymous):

okay let me see if this is right. 80*2^-56/28 = 80*2^-2 ?

OpenStudy (lgbasallote):

right. simplify that further

OpenStudy (anonymous):

=20

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

so back to one of the question. What does each value of the expression represent? would this mean that the equations i did equals to ? represent grams?

OpenStudy (lgbasallote):

i believe it means the answer to each one

OpenStudy (anonymous):

oh, okay.. lol thank you.

OpenStudy (lgbasallote):

welcome